OpenStudy (anonymous):

hi i need help with a word problem how would you write if the quotient of 13 and a number is added to six, the result is equal to 5 divided by the same number squared?

OpenStudy (dumbcow):

\[6+\frac{13}{n} = \frac{5}{n^{2}}\]

OpenStudy (dumbcow):

yielding a quadratic equation 6n^2 +13n -5 = 0 6n^2 -2n +15n -5 = 0 2n(3n-1) +5(3n-1) = 0 (3n-1)(2n+5) = 0 n = 1/3 n = -5/2

OpenStudy (anonymous):

thank you so much i just have a question why did you put a - sign after 13/n? wouldn't it be an = sign? and there are two answers? 1/3 and -5/2?

OpenStudy (dumbcow):

yes it would be an = sign, from my view it looks like an = sign and yes 2 solutions